OpenStudy (anonymous):

which of the following is a factor of this equation: 4-(x + Y)^2 -(x + y)^2 x + y 2 - x + y 2 + x + y 4 + x + Y

OpenStudy (anonymous):

You need to be careful with your brackets...

OpenStudy (anonymous):

\[4-(x+y)^2=(2-(x+y))(2+(x+y))\]

OpenStudy (anonymous):

Difference of two squares where one of your squares is 4, the other is (x+y)^2

OpenStudy (anonymous):

ohhhh i see, it's a multiple choice and i just couldn't figure out where they were getting the right answer

OpenStudy (anonymous):

so the correcy answer is 2 - x + y, why can't it be 2 + x + Y based on what you said?

OpenStudy (anonymous):

It goes back to what I was saying about the brackets. You have to do some rearranging and put it in the appropriate form. What you could do is just expand out the factors that I gave you and then match them up. So\[(2-(x+y))(2+(x+y))=(2-x-y)(2+x+y)\]

OpenStudy (anonymous):

When you look at your list, you only have 2 + x + y available to you.

OpenStudy (anonymous):

oh, because of the negative sign, that where the bracket comes into play. it makes sense now, thank you very much

OpenStudy (anonymous):

no worries. fan me if you haven't :) points hungry

OpenStudy (anonymous):

will do...first time on the site, how do i fan you?

OpenStudy (anonymous):

you just check a link that says "Become a fan" and leave it

OpenStudy (anonymous):

it's next to my name...and it's blue